When a glass prism of refracting angle 60o
is immersed in a liquid, its angle of minimum
deviation is 30o
. The critical angle of glass with respect to the liquid medium is

a) 45o
b) 30o
c) 60o
d) 55

Grade:

1 Answers

Roman
6 years ago
Refractive index of glass w.r.t liquid= sin((A+$)/2)/sin(A/2) = sin((60+30)/2/sin(60/2)= sin45/sin30= (1/root2)/(1/2)= root2 andRefractive index of glass w.r.t liquid= 1/sinc= root2 => sinc=1/root2 => c=45°Therefore option a) 45°